OpenStudy (anonymous):

Find an equation of a parabola that has the given numbers as solutions: (2-3i), (2+3i) I know it equals 13 but how do you find the equation?

OpenStudy (anonymous):

your job it to multiply out \[(x-(2-3i))(x-(2+3i))\] and that is not as hard as it looks. you already have the constant, it is \[2^2+3^2=13\]

OpenStudy (anonymous):

so all you need it the "middle term" which is \[-2x-2x=-4x\] and that is it. your answer is \[x^2-4x+13=0\] as the initial equation